There are 10 identified points on a number line. How many possible rays can
be drawn using the given points?

Found 2 solutions by amarjeeth123, ikleyn:
Answer by amarjeeth123(569)

We need two points to form a ray.
This is a combination question.
Number of possible rays=C(10,2)=(10!)/(8!)(2!)=45
Answer=45 rays.

Answer by ikleyn(52777)

.


            Tutor @amarjeeth is not right,  and his answer is not correct.



If to accept the conception,  that a ray is determined by a pair of points,
then these pairs must be considered as  ORDERED  (an order does matter),

and the number of different pairs and different rays will be  10*9 = 90  (not  45  (!) ).


In this problem,  we work with  PERMUTATIONS  of  10  points, taken  2  at a time,  and  NOT  with the  COMBINATIONS  (!).


Again,  the order of points in pairs  DOES  MATTER  in this problem (!).


-------------

Solved  (correctly).